User Tools

Site Tools


operations:agenda2016_10_13

Chair/Presenting: Ellen/

In Attendance: Nick, Liz, Bryn, Jamie, Lucia

Apologies: Jonny, Ross, Arwin, Tiege, Lucas, Jim

Absent: Jesse

___

Agenda:

General Issues

NOTE: This is a compulsory meeting for observers, you need a good reason not to come - if you have lectures we can change the time

  • LBA next week - keep skype open and check the LBA chat, let atnf people know when you come on shift
  • Schedule is at the link below, instructions are on the ra wiki

http://www.atnf.csiro.au/vlbi/dokuwiki/doku.php/lbaops/lbaoct2016/start

  • Is the power still down at Mt Pleasant, do we have an ETA? - back on now, everything is good, anemometer is still iffy, might not stow in high winds (Brett is working on now)
  • Caltsys at Ho not running means overflowing (useless) tsys values. DO THE CHECKLIST PROPERLY EVERY 2 HOURS.
    • Cal was setup but must have dropped out
  • Does Arwin's windstow script work? Has anyone had any problems?
  • Is the new flogit script working now? Before it was giving log file .prc extension - seems to be working
  • Warren sent out an email re module allocation - check spreadsheet during experiment set up (rapids shouldn't go on modules with anything else, except each other)

http://auscope.phys.utas.edu.au/opswiki/doku.php?id=operations:log

  • Make sure you check the end of experiment email before sending it - especially how much data is on each module if multiple are used, this should be recorded in the handover notes when the banks swap over
  • Check the summary file for each telescope when choosing a module to use, if at all possible, make sure it's all on one module - the experiment shouldn't fill up the module 2 hours before the end. This is less of an issue for Hobart for some experiments. - Next time, swap modules, don't just lose those 2 hours (can be copied off very quickly if the module needs to go somewhere else)

Issues from Handover notes

  • AUA013 - during setsup ifc power levels were below target, 27K when expecting 32k (Arwin?) - do we know what happend with this, did it fix itself
  • If you spot a problem either fix it or keep an eye on it. Don't just report it and do nothing, even if it's just saying the on-call person said it couldn't be fixed.
    • 11:20:00 UT Temporarily got a GPIB time-out when maserdelay called, set off alarm for Yg and Hb. Counters were fine and within 0.3 us difference. Monitoring. (Lucas)
  • If an experiment starts late, put the reason and the first full scan in the handover notes. Observers don't get the start messages so the information needs to be in the hand over notes. You also can't track the disc position very well if you don't know how much was lost at the beginning…
    • start messages are to be sent out even if a telescope can't participate, if they are scheduled to be in the experiment, you need to let the correlator know they shouldn't expect data
    • more details in start and end messages - ie missed scans are a problem to be noted
  • R4760 - 08:00UT Is S-band right polarisation attenuation now supposed to be set at 6 dB, has been noted as either 0 or 6 in the setup instructions but not with much confidence, used to be 7 I think (Ross).
    • Brett to have a look at power levels
  • Dynamic scheduling seems to have quite regular slewing during scan errors
    • Please keep handover notes so Lucia can improve them

New monitors are on the way :)

There is a script (line of code) to determine previous disc positions - written on paper “Dave's Legacy”

Bryn is chairing and presenting next fortnight (27th).

IT Resources: complain to them about internet dropping out for whole day in whole building (re: 1st Oct)

Jamie to do: fix flogit for Ho to take the right log file rather than vela's

Jim to look at: weather data a Yg regularly has huge gaps (half days)

/home/www/auscope/opswiki/data/pages/operations/agenda2016_10_13.txt · Last modified: 2016/10/13 02:43 by Ellen Manning